A rock, which weighs 1400 N in air, is dropped into fresh water (998 kg/m3). You measure that it accelerates downward at a rate of 6.25 m/s2. What is the volume of the rock?
Answer #1

net firce acting on rock

ma = W - F

(1400/9.8} * 6.25 = 1400 - 1000* V * 9.8

V = 0.0517 m^3

==========

comment before rate in case any doubt, will reply for sure.. goodluck